Solve.
8x-(3x+6)=4

Answer by nerdybill(7384)

8x-(3x+6)=4
distribute negative sign to each term inside parenthesis:
8x-3x-6=4
combine like-terms:
5x-6=4
add 6 to both sides:
5x = 10
divide both sides by 5:
x = 2